OpenStudy (anonymous):

x+y+z=9 w+y+z=5 w+x=2 What is the average of w, x, y, and z above?

OpenStudy (anonymous):

x+y+z=9......(i) w+y+z=5......(ii) w+x=2......(iii) from (iii) x = 2-w substituting in (i) 2-w+y+z =9 ; y+z-w =7....(iv) add (ii) and (iv) (w+y+z)+( y+z-w) =7+5 = 12 w-w +y+y+z+z = 12 2y+2z =12 y+z =6 ....(v) now add (iii) and (v) together (w+x) +(y+z) =6+2 = 8 w+x+y+z =8.....(vi) now from (vi) you can get the average which is simply 8/4 = 2